OpenStudy (anonymous):

Need step-by step help on this one...not just the answer...need to know how to get there please log6 3+ log6 8 - log6 4

OpenStudy (anonymous):

hello Jenn we shall work on this

OpenStudy (anonymous):

First observe that every log has same base..

OpenStudy (anonymous):

i Need steps please... There are all log6 - are these considered like terms?

OpenStudy (anonymous):

yes.. log to the same base are considered as like ones

OpenStudy (anonymous):

Good thought, but because the argument of the log is different they are each different numbers.

OpenStudy (anonymous):

how do I proceed?

OpenStudy (anonymous):

Logk a + logk b = logk ab

OpenStudy (anonymous):

They can be combined using the addition property of logs

OpenStudy (anonymous):

which is actually multiplication?

OpenStudy (anonymous):

If logs are combined by addition then multiply the inner terms else if subtraction then divide

OpenStudy (anonymous):

As helping has explained.\[log_b(a) + log_b(c) = log_b(a\cdot c)\]

OpenStudy (anonymous):

do I work with the first 2 tems, and then the last 2 terms? Multiply 3*8, then get log6 24, then divide by the log 6 4?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

You don't divide by the log, you divide the argument

OpenStudy (anonymous):

Just divide the inner terms

OpenStudy (anonymous):

So then I would get log6 6?

OpenStudy (anonymous):

like logk a - logk b = logk a/b

OpenStudy (anonymous):

\[log_b(a) + log_b(c) - log_b(d) = log_b(\frac{a\cdot c}{d})\]

OpenStudy (anonymous):

That's right. log6(6)

OpenStudy (anonymous):

Which is just 1 right?

OpenStudy (anonymous):

I can work with formulas, but the instructions that I have, don't offer the ones that the homework questions are asking me to solve! Curious, the writers of these things get paid, why?

OpenStudy (anonymous):

You understand this about logs? \[\log_b(b) = 1\]

OpenStudy (anonymous):

Jenn,, logk k is 1, log of a number to same base is always 1

OpenStudy (anonymous):

if the log base and the number it has next to it are the same, then they equal 1?

OpenStudy (anonymous):

Yes absolutely

OpenStudy (anonymous):

MUCHAS GRACIAS! That is VERY helpful to know!

OpenStudy (anonymous):

Yes, because recall our exponential form: \[Let\ k = log_b(b) \implies b^k = b \implies k = 1 \implies log_b(b) = 1\]

OpenStudy (anonymous):

That again is super helpful. Just wrote it down for future reference!

OpenStudy (anonymous):

Also \[log_b(1) =0\]

OpenStudy (anonymous):

Because if we do the same thing: \[Let \ k = log_b(1)\]\[\implies b^k = 1 = b^0\]\[\implies k = 0\]\[\implies log_b(1) = 0\]

OpenStudy (anonymous):

you are really a good student to work with,,,

OpenStudy (anonymous):

I just posted another one that need your attention please!